# Faster Cascades via Speculative Decoding

### v1

@@ Abstract
Cascades and speculative decoding are two common approaches to improving language models’ inference efficiency. Both approaches involve interleaving models of different sizes, but via fundamentally distinct mechanisms: cascades employ a deferral rule that invokes the larger model only for “hard” inputs, while speculative decoding uses speculative execution to primarily invoke the larger model in parallel verification mode. These mechanisms offer different benefits: empirically, cascades are often capable of yielding better quality than even the larger model, while theoretically, speculative decoding offers a guarantee of quality-neutrality. In this paper, we leverage the best of both these approaches by designing new speculative cascading techniques that implement their deferral rule through speculative execution. We characterize the optimal deferral rule for our speculative cascades, and employ a plug-in approximation to the optimal rule. Through experiments with T5 models on benchmark language tasks, we show that the proposed approach yields better cost-quality trade-offs than cascading and speculative decoding baselines.
@@ 1 Introduction
Large language models (LLMs) have demonstrated significant advances in quality on a range of natural language processing tasks [33, 34, 5, 2, 11, 48, 12, 41, 1, 43, 42], at the cost of a significant increase in inference latency. This has sparked a growing body of literature on reducing LMs’ inference costs without (overly) compromising on quality [15, 32, 36, 26, 7, 37, 40]. One such line of work involves constructing a family of models of various sizes (e.g., a small and large model), and suitably orchestrating amongst them to make a prediction. Two canonical instantiations of this strategy are model cascading [47, 28, 45, 23, 14, 8, 18, 13] and speculative decoding [39, 7, 26, 40, 49].
@@
Owing to their different mechnanisms, both methods have complementary strengths. Cascades seek to output distributions that have the best quality for a given cost budget, and are empirically observed to often yield better accuracies than even the individual models they are constructed with [22, 24]3). By contrast, speculative decoding is theoretically guaranteed to match the output distribution (or a close approximation thereof [44]), and are practically observed to provide impressive speed-ups [39, 7, 26, 40]. Given the complementary nature of these two approaches, a natural question that arises is whether we can leverage the best of both techniques.
In this paper, we do so by designing new techniques for two-model cascades that implement their deferral rule in a speculative manner: we have the smaller model generate drafts auto-regressively, and the larger model execute in parallel on the drafts to decide whether or not to defer on themWe show that this speculative cascading approach yields better cost-quality trade-offs than both standard cascades and speculative decoding. In detail, we make the following contributions:
@@
    We introduce a general recipe for speculative execution, where we seek to mimic a general target distribution that interleaves the drafter’s and verifier’s output distributions. Lossy speculative sampling [44] is a special case of this recipe for a particular target distribution (§4.1).
@@
    We show how common cascading deferral rules, such as Chow’s rule [10] and confidence-difference thresholding [22], can be implemented speculatively by plugging in the corresponding target distribution into the proposed framework. We refer to these as speculative cascades4.2).
@@
    We characterize the theoretically optimal deferral rule for a speculative cascade, and design a speculative cascading technique that implements a plug-in estimate to the optimal rule (Lem. 4).
@@
    Through experiments on benchmark language tasks, we show that speculative cascades constructed from T5 models [35] of different sizes are able to provide better cost-quality trade-offs than their sequential cascade and speculative decoding counterparts (§6).
@@ 2 A Tale of Two Efficient LM Inference Strategies
Let $\mathscr{V}$ denote a finite vocabulary of tokens, with $\mathscr{V}^{*}$ the set of all sequences generated by this vocabulary. Let $\Delta_{\mathscr{V}}$ denote the set of all probability distributions over tokens in $\mathscr{V}$. Given an arbitrary length sequence $x=x_{1}x_{2}\ldots x_{L}\in\mathscr{V}^{*}$ and index $i\leq L$, denote by $x_{<i}=x_{1}x_{2}\ldots x_{i-1}$. A language model (LM) is a probability distribution over $\mathscr{V}^{*}$. Let $\mathbb{P}$ denote the ground-truth probability distribution over $\mathscr{V}^{*}$. This could be, for example, a distribution over prompt-response pairs that the LM may encounter during deployment, or a distribution of sequences used to pre-train the LM. We will measure the quality of an LM based on how closely it mimics $\mathbb{P}$.
@@
Suppose we are provided two LMs $q$ and $p$, where $p$ is the larger model. Our goal is to design an inference strategy that selectively invokes $q$ and $p$ to trade-off between quality and latency (which may be approximated by the fraction of times that $p$ is invoked). We will denote by $q(x_{t}|x_{<t})$ the probability $q$ associates to token $x_{t}\in\mathscr{V}$ given prefix $x_{<t}\in\mathscr{V}^{t-1}$, and by $p(x_{t}|x_{<t})$ the same distribution from model $p$. Whenever it is clear from context, we will hide the conditioning on prefix $x_{<t}$, and use the shorthand $p_{t}(\cdot)$ for $p(\cdot|x_{<t})$ and $q_{t}(\cdot)$ for $q(\cdot|x_{<t})$.
Cascades.
@@
Cascades are an effective strategy to trade-off cost and quality by having the smaller model $q$ handle the “easy” samples, and the larger model $p$ handle the “difficult” samples [18, 50]. A common approach to cascading is confidence-based thresholding or Chow’s rule [10, 22], where we first run $q$ on the given input, and defer to $p$ when $q$’s confidence for its generated response is sufficiently low. This strategy is typically implemented at the sequence-level, where for a given prefix $x_{<m}$ we invoke $q$ to generate a complete response $x_{m}\ldots x_{m+n}$. We evaluate the predicted probability from $q$ for this response, and check whether it falls below a threshold $\alpha\in[0,1]$:
$q(x_{m}\ldots x_{m+n}\,|\,x_{<m})<1-\alpha.$ (1)
@@
If the above holds, we defer to $p$ to generate a new response; otherwise, we retain $q$’s response. One may then tune the threshold to achieve a desired cost-quality trade-off. The literature also offers variants of Chow’s rule that use a more nuanced aggregation of per-token uncertainties [18].
Speculative decoding.
@@
Speculative decoding is an alternate inference strategy that applies token-level interleaving between $q$ and $p$, resulting in provably matching the larger model quality at a reduced inference cost [39, 26]. Given a prefix $x_{<t}$, we sample $\gamma$ draft tokens $x_{t},\ldots,x_{t+\gamma-1}$ auto-regressively from $q$, and run $p$ in parallel on the $\gamma$ prefixes $x_{<t},\ldots,x_{<t+\gamma-1}$, and verify if the generated tokens can be accepted. We then rollback to the first token $t+j^{*}$ that was rejected, replace $x_{t+j^{*}}$with a new token, and repeat the drafting-verification process with $x_{<t+j^{*}+1}$ as the new prefix.
@@
During the verification stage, a draft token $x_{t+j}$ generated by $q$ is accepted with probability $\min\left(1,\frac{p_{t+j}(x_{t+j})}{q_{t+j}(x_{t+j})}\right)$ and rejected otherwise, where recall the shorthand $q_{t+j}(\cdot)=q(\cdot|x_{<t+j})$ and $p_{t+j}(\cdot)=p(\cdot|x_{<t+j})$. A rejected token is then replaced by a new token sampled from a modified distribution $\operatorname{norm}\left(\max\left\{0,\,p_{t+j}(\cdot)-q_{t+j}(\cdot)\right\}\right),$ where $\text{norm}(\cdot)$ denotes normalization to sum to 1. This sampling process is provably equivalent to sampling $\gamma$ tokens auto-regressively from $p$ for prefix $x_{<t}$ [26]. We summarize this speculative sampling procedure in Algorithm 1. Each invocation of this algorithm generates at most $\gamma+1$ next tokens for a given prefix $x_{<t}$. One may run this algorithm multiple times to generate a complete output sequence.
@@
In practice, one may employ a lossy variant [44] of the above sampling that allows some deviation from verifier’s distribution $p$. In this case, a draft token $x_{t+j}$ is accepted with probability $\min\left(1,\frac{p_{t+j}(x_{t+j})}{(1-\alpha)\cdot q_{t+j}(x_{t+j})}\right)$, where $\alpha\in[0,1)$ is a strictness parameter, with higher values indicating greater deviation from $p$. A rejected token may then be replaced by a token sampled from the residual distribution $\operatorname{norm}\left(\max\left\{0,\,\frac{1}{\beta}\cdot p_{t+j}(\cdot)-q_{t+j}(\cdot)\right\}\right),$ where $\beta\geq 1-\alpha$ is a parameter that depends on $\alpha$, $q$ and $p$. A common heuristic is to simply set $\beta=1$ [52].
@@
SpecDecoding [26] - $p(x)$ Speculative
Lossy SpecDecoding [44] - $\max\{\min\{q(x),\frac{p(x)}{1-\alpha}\},\frac{p(x)}{\beta}\}$ Speculative
BiLD* [24] $\mathbf{1}\big{(}\,D(q,p)>\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
TokenCascade [Chow] [10] $\mathbf{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Sequential
Oracle [Diff] [22] $\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Oracle
SpecCascade [Chow] $\mathbf{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [Diff] $\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [OPT] $\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\cdot D_{\textrm{TV}}(p,q)\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
Table 1: Target distributions associated with different inference algorithms, where $\alpha$ is a free parameter and $\beta\geq 1-\alpha$ is a parameter dependent on $\alpha$, $q$ and $p$. The last column indicates whether the execution is sequential (Algorithm 2), via an oracle (Algorithm 3), or speculative (Algorithm 5) [26]. The third row presents a variant of the BiLD algorithm of [24], where $D(q,p)$ is a measure of discrepancy between $q$ and $p$; the original algorithm differs in the use of a deterministic speculative decoding procedure with a dynamic draft window (see §5). See (6) for more details on $\delta$.
@@ 3.1 Warm-up: Token-level cascades
It is straight-forward to extend the sequence-level Chow’s rule from §2 to form a token-level cascade between $q$ and $p$. For a prefix $x_{<t}$, we first compute the smaller model’s distribution $q(\cdot|x_{<t})$, and check whether $\max_{v\in\mathscr{V}}\,q(v|x_{<t})$ is below a pre-chosen threshold. if so, we evaluate $p(\cdot|x_{<t})$, and sample $x_{t}\sim p(\cdot|x_{<t})$; otherwise, we sample $x_{t}\sim q(\cdot|x_{<t})$.
@@
${r}_{\rm\tt Chow}(x_{<t})=1~{}\iff~{}\max_{v\in\mathscr{V}}\,q(v|x_{<t})<1-\alpha,$ (2)
where $\alpha$ is a cost parameter; the higher the value, the lower is the frequency of deferral to $p$. One may also use a confidence measure different from the maximum probability, such as the entropy of the small model’s probability distribution. We elaborate in §B that the choice of confidence measure would depend on the evaluation metric of interest; (2) is typically prescribed when the cascade’s quality is evaluated in terms of its accuracy against the ground-truth distribution on individual tokens, whereas entropy is prescribed when the metric of interest is the cross-entropy loss.

### v2

@@ Abstract
Cascades and speculative decoding are two common approaches to improving language models’ inference efficiency. Both approaches involve interleaving models of different sizes, but via fundamentally distinct mechanisms: cascades employ a deferral rule that invokes the larger model only for “hard” inputs, while speculative decoding uses speculative execution to primarily invoke the larger model in parallel verification mode. These mechanisms offer different benefits: empirically, cascades offer better cost-quality trade-offs, often even outperforming the large model, while theoretically, speculative decoding offers a guarantee of quality-neutrality. In this paper, we leverage the best of both these approaches by designing new speculative cascading techniques that implement their deferral rule through speculative execution. We characterize the optimal deferral rule for our speculative cascades, and employ a plug-in approximation to the optimal rule. Experiments with Gemma and T5 models on a range of language benchmarks show that our approach yields better cost-quality trade-offs than cascading and speculative decoding baselines.
@@ 1 Introduction
Large language models (LLMs) have yielded significant advances in quality on a range of natural language processing tasks (Radford et al., 2018; Raffel et al., 2020a; Brown et al., 2020; Black et al., 2022; Chowdhery et al., 2022; Wei et al., 2022; Chung et al., 2022; Tay et al., 2023; Anil & et al., 2023; Touvron et al., 2023; Team et al., 2023), at the cost of an increase in inference latency. This has sparked a growing body of literature on reducing LMs’ inference costs without (overly) compromising on quality (Elbayad et al., 2020; Pope et al., 2022; Schuster et al., 2022; Leviathan et al., 2023; Chen et al., 2023a; Sheng et al., 2023; Sun et al., 2024). One such line of work involves constructing a family of models of various sizes (e.g., a small and large model), and suitably orchestrating amongst them to make a prediction. Two canonical instantiations of this strategy are model cascading (Wang et al., 2020; Mamou et al., 2022; Varshney & Baral, 2022; Khalili et al., 2022; Dohan et al., 2022; Chen et al., 2023b; Gupta et al., 2024; Ding et al., 2024) and speculative decoding (Stern et al., 2018; Chen et al., 2023a; Leviathan et al., 2023; Sun et al., 2024; Li et al., 2024a; Xia et al., 2024a).
@@
Owing to their different mechanisms, both methods have complementary strengths. Cascades seek to output distributions that have the best quality for a given cost budget, and potentially provide better cost-quality trade-offs, sometimes even yielding better accuracies than the individual models they are constructed with (Jitkrittum et al., 2023; Kim et al., 2023)3). By contrast, speculative decoding is theoretically guaranteed to match the output distribution (or a close approximation thereof (Tran-Thien, 2023)), and are practically observed to provide impressive speed-ups (Stern et al., 2018; Chen et al., 2023a; Leviathan et al., 2023; Sun et al., 2024). Given the complementary nature of these two approaches, a natural question arises: can we leverage the best of both techniques?
In this paper, we do so by designing new techniques for two-model cascades that implement their deferral rule in a speculative manner: we have the smaller model generate drafts auto-regressively, and the larger model execute in parallel on the drafts to decide whether or not to defer on them. We show that this speculative cascading approach yields better cost-quality trade-offs than both standard cascades and speculative decoding. In detail, we make the following contributions:
@@
    We introduce a general recipe for speculative execution, where we seek to mimic a general target distribution that interleaves the drafter’s and verifier’s distributions. Lossy speculative sampling (Tran-Thien, 2023) is a special case of this recipe for a particular target distribution (§4.1).
@@
    We show how common cascading rules, such as Chow’s rule (Chow, 1970) and confidence-difference thresholding (Jitkrittum et al., 2023), can be implemented speculatively by plugging in their target distribution into our framework. We refer to these as speculative cascades4.2).
@@
    We characterize the theoretically optimal deferral rule for a speculative cascade, and design a speculative cascading technique that implements a plug-in estimate to the optimal rule (§4.3, Lemma 4, Table 1). We also present token-specific variants of our deferral rules (§4.4).
@@
    Through experiments with Gemma (Team et al., 2024) and T5 models (Raffel et al., 2020b) on a range of benchmark language tasks including summarization, translation, reasoning, coding and QA, we show that speculative cascades are able to provide better cost-quality trade-offs than their sequential cascade and speculative decoding counterparts (§6).
@@ 2 A Tale of Two Efficient LM Inference Strategies
Let $\mathscr{V}$ denote a finite vocabulary of tokens, with $\mathscr{V}^{*}$ denoting the set of all sequences generated by this vocabulary. Let $\Delta_{\mathscr{V}}$ denote the set of all probability distributions over tokens in $\mathscr{V}$. Given an arbitrary length sequence $x=x_{1}x_{2}\ldots x_{L}\in\mathscr{V}^{*}$ and index $i\leq L$, denote by $x_{<i}=x_{1}x_{2}\ldots x_{i-1}$. A language model (LM) is a probability distribution over $\mathscr{V}^{*}$. Let $\mathbb{P}$ denote the ground-truth probability distribution over $\mathscr{V}^{*}$. This could be, for example, a distribution over prompt-response pairs that the LM may encounter during deployment, or a distribution of sequences used to pre-train the LM. We will measure the quality of an LM based on how closely it mimics $\mathbb{P}$.
@@
Suppose we are provided two LMs $q$ and $p$, where $p$ is the larger (more expensive) model. Our goal is to design an inference strategy that selectively invokes $q$ and $p$ to trade-off between quality and latency (which may be approximated by the fraction of times that $p$ is invoked). We will denote by $q(x_{t}|x_{<t})$ the probability $q$ associates to token $x_{t}\in\mathscr{V}$ given prefix $x_{<t}\in\mathscr{V}^{t-1}$, and by $p(x_{t}|x_{<t})$ the same distribution from model $p$. Whenever it is clear from context, we will hide the conditioning on prefix $x_{<t}$, and use the shorthand $p_{t}(\cdot)$ for $p(\cdot|x_{<t})$ and $q_{t}(\cdot)$ for $q(\cdot|x_{<t})$.
@@
Cascades are an effective strategy to trade-off cost and quality by having the smaller model $q$ handle the “easy” samples, and the larger model $p$ handle the “hard” ones (Gupta et al., 2024; Yue et al., 2024). A common cascading approach is confidence thresholding or Chow’s rule (Chow, 1970; Jitkrittum et al., 2023), where we first run $q$ on the input, and defer to $p$ when $q$’s confidence for its generated response is sufficiently low. This strategy is typically implemented at the sequence-level, where for a given prefix $x_{<m}$ we invoke $q$ to generate a complete response $x_{m}\ldots x_{m+n}$. We evaluate $q$’s predicted probability for the response, and check whether it falls below a threshold $\alpha\in[0,1]$:
$\textstyle q(x_{m}\ldots x_{m+n}\,|\,x_{<m})<1-\alpha.$ (1)
@@
If the above holds, we defer to $p$ to generate a new response; otherwise, we retain $q$’s response. One may tune the threshold to achieve a desired cost-quality trade-off. The literature also offers variants of Chow’s rule that use a more nuanced aggregation of per-token uncertainties (Gupta et al., 2024).
@@
Speculative decoding is an alternate strategy that applies token-level interleaving between $q$ and $p$, resulting in provably matching the larger model quality at a reduced inference cost (Stern et al., 2018; Leviathan et al., 2023). Given a prefix $x_{<t}$, we draft $\gamma$ draft tokens $x_{t},\ldots,x_{t+\gamma-1}$ via auto-regressive sampling from $q$, and verify if these tokens can be accepted by running $p$ in parallel on the $\gamma$ prefixes $x_{<t},\ldots,x_{<t+\gamma-1}$. We then rollback to the first rejected token $t+j^{*}$ (where $j^{*}\in\{0,1,\ldots,\gamma-1\}$), replace $x_{t+j^{*}}$ with a new token, and repeat the process with prefix $x_{<t+j^{*}+1}$.
@@
During the verification stage, a draft token $x_{t+j}$ generated by $q$ is accepted with probability $\min\left(1,\frac{p_{t+j}(x_{t+j})}{q_{t+j}(x_{t+j})}\right)$ and rejected otherwise, recalling the shorthand $q_{t+j}(\cdot)=q(\cdot|x_{<t+j})$ and $p_{t+j}(\cdot)=p(\cdot|x_{<t+j})$. A rejected token is then replaced by a new token sampled from a modified distribution $\operatorname{norm}\left(\max\left\{0,\,p_{t+j}(\cdot)-q_{t+j}(\cdot)\right\}\right),$ where $\text{norm}(\cdot)$ denotes normalization to sum to 1. This sampling process is provably equivalent to sampling $\gamma$ tokens auto-regressively from $p$ for prefix $x_{<t}$ (Leviathan et al., 2023). We summarize this speculative sampling procedure in Algorithm 1. Each invocation of this algorithm generates at most $\gamma+1$ next tokens (and at least one) color=RoyalBlue!35!white]and at least $1$? Hari: done! for a given prefix $x_{<t}$. One may run this algorithm multiple times to generate a complete output sequence.
@@
In practice, one may employ a lossy variant (Tran-Thien, 2023) of the above sampling that allows some deviation from verifier’s distribution $p$. In this case, a draft token $x_{t+j}$ is accepted with probability $\min\left(1,\frac{p_{t+j}(x_{t+j})}{(1-\alpha)\cdot q_{t+j}(x_{t+j})}\right)$, where $\alpha\in[0,1)$ is a strictness parameter, with higher values indicating greater deviation from $p$. A rejected token may then be replaced by a token sampled from the residual distribution $\operatorname{norm}\left(\max\left\{0,\,\frac{1}{\beta}\cdot p_{t+j}(\cdot)-q_{t+j}(\cdot)\right\}\right),$ where $\beta\geq 1-\alpha$ is a parameter that depends on $\alpha$, $q$ and $p$. A common heuristic is to simply set $\beta=1$ (Zhou et al., 2024).
Table 1: Target distributions associated with different inference algorithms, where $\alpha$ is a free parameter and $\beta\geq 1-\alpha$ depends on $\alpha$, $q$ and $p$. The last column indicates whether the execution is sequential (Algorithm 2), via an oracle (Algorithm 3), or speculative (Algorithm 5) (Leviathan et al., 2023). See (6) for details on $\delta$. The third row presents a variant of the BiLD algorithm of Kim et al. (2023), where $D(q,p)$ is a measure of discrepancy between $q$ and $p$; the original algorithm differs in the use of a deterministic speculative decoding procedure with a dynamic draft window (see §5).
@@
SpecDecoding (Leviathan et al., 2023) - $p(x)$ Speculative
Lossy SpecDecoding (Tran-Thien, 2023) - $\max\{\min\{q(x),\frac{p(x)}{1-\alpha}\},\frac{p(x)}{\beta}\}$ Speculative
BiLD* (Kim et al., 2023) $\bm{1}\big{(}\,D(q,p)>\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
TokenCascade [Chow] (Chow, 1970) $\bm{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Sequential
Oracle [Diff] (Jitkrittum et al., 2023) $\bm{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Oracle
SpecCascade [Chow] $\bm{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [Diff] $\bm{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [OPT] $\bm{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\cdot D_{\textrm{TV}}(p,q)\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
@@ 3.1 Warm-up: Token-level cascades
It is straightforward to extend the sequence-level Chow’s rule from §2 to form a token-level cascade between $q$ and $p$. For a prefix $x_{<t}$, we first compute the smaller model’s distribution $q(\cdot|x_{<t})$, and check whether $\max_{v\in\mathscr{V}}\,q(v|x_{<t})$ is below a pre-chosen threshold. if so, we evaluate $p(\cdot|x_{<t})$, and sample $x_{t}\sim p(\cdot|x_{<t})$; otherwise, we sample $x_{t}\sim q(\cdot|x_{<t})$.
@@
$\textstyle{r}_{\rm\tt Chow}(x_{<t})=1\leavevmode\nobreak\ \iff\leavevmode\nobreak\ \max_{v\in\mathscr{V}}\,q(v|x_{<t})<1-\alpha,$ (2)
where $\alpha$ is a threshold parameter; the higher the value, the lower is the frequency of deferral to $p$. One may also use other confidence measures than the maximum probability, such as the entropy of the small model’s probability distribution. We elaborate in §B that the choice of confidence measure would depend on the evaluation metric of interest; Equation 2 is typically prescribed when the cascade’s quality is evaluated in terms of its accuracy against the ground-truth distribution on individual tokens, whereas entropy is prescribed when the metric of interest is the cross-entropy loss.